OpenStudy (anonymous):

ln(x-4)-ln(x)=2 Solve for x. Help Please

OpenStudy (anonymous):

\[\ln \left( x-4 \right)-\ln \left( x \right)=2\]

OpenStudy (anonymous):

No solution